OpenStudy (anonymous):

What is the sum of the sequence: 1+3+5+7+...1001

OpenStudy (anonymous):

I have 1004003 is this correct?

OpenStudy (anonymous):

the sum of consecutive odd numbers is the square of how many terms you have \[1=1\] \[1+3=4=2^2\] \[1+3+5=9=3^2\] \[1+3+5+7=16=4^2\] \[1+3+5+7+9=25=5^2\] so the question is, how many odd numbers do you have?

OpenStudy (anonymous):

hint, odd numbers look like \[2n-1\] so set \[2n-1=1001\] and solve for n. the square your result

OpenStudy (anonymous):

your answer is not correct, but i am sure you can get it

OpenStudy (anonymous):

no clue :( thanks anyhow